OpenStudy (anonymous):

if repetitions are not allowed, how many four digit numbers that a) are even b) greater than 7000 can be formed from the digits 1,2, ...,8,9

OpenStudy (mathmate):

Case 1: the number starts with a 7 or a 9 There are 2 choices for the first digit, 5 choices for the last, (even) and 8 for the second and 7 for the third for a total of 2*5*8*7=560 numbers Case 2: the number starts with an 8 there is one choice for the first digit, 4 choices for the last, and 8 for the second and 7 for the third, for a total of 1*4*8*7=224 numbers Sum of cases 1 & 2 = 560+224 = 784 even numbers between 7000 and 9999

OpenStudy (anonymous):

oh the answer in the solution booklet says that it a) is 1344; any ideas how they got this?

OpenStudy (mathmate):

Was the answer 1344 for even numbers between 7000 and 9999 without repetition?

OpenStudy (anonymous):

oh it didn't say, so i'm not sure

OpenStudy (mathmate):

My bad, the digits do not include zero, and the requirements are not combined for a and b. I`ll start over: a. Even numbers between 7000 and 9999 without repetition 4 digits for the last 8 digits for the third 7 digits for the second 6 digits for the first That makes 4*8*7*6=1344 numbers.

OpenStudy (mathmate):

b. 4 digit numbers greater than 7000: 3 choices for the first digit (7,8 or 9) 8 choices for the second 7 choices for third and 6 digits for the fourth, for a total of 3*8*7*6 = 1008

OpenStudy (anonymous):

OH I SEEE! HEHEHE THANK YOU

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

would it be ok if you helped me with b) thanks mathmate!

OpenStudy (mathmate):

Done! Check post following (a).

OpenStudy (anonymous):

oh because the ans in the solution book is 1008

OpenStudy (mathmate):

That's what I gave for (b). You may have missed the two posts send one after the other. The second one was sent "5 minutes ago".

OpenStudy (anonymous):

ahhhhhhh MY APOLOGIES!!!!!!!!!!!!!!!!!

OpenStudy (mathmate):

No problem at all! :)

OpenStudy (anonymous):

@mathmate, are you sure you are counting even numbers between 7000 and 9999 without repetition in as: >>4 digits for the last >>8 digits for the third >>7 digits for the second >>6 digits for the first >>That makes 4*8*7*6=1344 numbers.

OpenStudy (mathmate):

Not really, what I got was "Sum of cases 1 & 2 = 560+224 = 784 even numbers between 7000 and 9999" was not what the question asked for. It was divided into two cases where the number starts with 7 or 9 (case 1) or starts with 8 (case 2).